Keystone XL pipeline-Alberta to the Gulf of Mexico TransCanada says building the 57 bilion pieline Explain how TransCanadas investment of S7 bilion could create spin-ofl jobs and pump $20 billion into the U.S economy TransCanadas investment of $7 bilion could create spin-off jobs and pump $20 billion into the US economy because A. new employees spend their incomes on consumption goods and services, creating spin-offjobs and more production would put 20.000 Americans directly to work during the construction phase and add an B. potential GDP will increase by $7 billion, which induces an increase in aggregate demand O C. TransCanada estimates that the multiplier is 3 5 ○ D. TransCanada estimates that mapier is lessthan Clck to select your answer Type here to search

maricans frecity to work during the corsltuctfion phase and add an espected 118.000 spihi offjsls, The project would also pump $20 tbillon into the U.S. economy n off jols and pump $20 biion into the U S economy d pump $20 bilion ins the US economy because and services, oreating spin-off jobs and more prodaction increase in aggregate demand
Answer #1

Transcanada's investment of $7 billion could create spin off jobs and pumps 20 billion into the US economy because  new employees spend their income on consumption of goods and services, creating spin off jobs and more production

Ans is a) new employees spend their income on consumption of goods and services, creating spin off jobs and more production

Estimated multiplier ;

7*m = 20

m = 20/7 = 2.86( approximately)

so option c) and d) is false.

Option b) is also incorrect as aggregate demand does not increase if potential GDP increases.

**next question is blurred and also part of it is not in the picture ( please provide a proper picture for the next question)